Arcadiouse (arcadiouse):

2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=??

OpenStudy (will.h):

beautiful

Awolflover1 (awolflover1):

well there are 7 wolves and 8 coyotes how many are there in total ?

Arcadiouse (arcadiouse):

98 2+3=2*[3+(2-1)]=8 3+7=3*[7+(3-1)]=27 4+5=4*[5+(4-1)]=32 5+8=5*[8+(5-1)]=60 6+7=6*[7+(6-1)]=72 therefore 7+8=7*[8+(7-1)]=98 x+y=x[y+(x-1)]=x^2+xy-x

Awolflover1 (awolflover1):

lol wow i see xD

OpenStudy (will.h):

well. it's still beautiful when i say beautiful i mean awful

OpenStudy (will.h):

i noticed that the values to the left increases by one. uh that was hard to know what for.

Arcadiouse (arcadiouse):

Thanks lol. I didn't think anyone would figure it out.